Carla and benicio work in a men's clothing store. They earn commission from each suit and each pair of shoes they sell. For selling 3 suits and 1 pair of shoes
shoes, Carla has earned $47 in commission. For selling 7 suits and 2 pairs of shoes, benicio has earned $107 in commission. How much do the salespeople earn for the sale of a suit? How much do they earn for the sale of a pair pair of shoes?

Answer by TimothyLamb(4379)

x = dollars commission for one suit
y = dollars commission for one pair shoes
---
3x + 1y = 47
7x + 2y = 107
---
put the system of linear equations into standard form
---
3x + 1y = 47
7x + 2y = 107
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 13
y = 8
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---